OpenStudy (anonymous):

Differentiate with repect to x, where y= 3x

OpenStudy (aravindg):

dy/dx= d/dx(3x) 3 is a constant so take outside

OpenStudy (aravindg):

dy/dx=3 d/dx(x)

OpenStudy (aravindg):

now just apply derivative of \[\large x^n=nx^{n-1}\] Here n=1

OpenStudy (aravindg):

got it ?

OpenStudy (anonymous):

Wait I you draw it out please? ckets and stuff. Can you draw it out please?

OpenStudy (aravindg):

ckets?

OpenStudy (anonymous):

brackets sorry i always DISCONNECT

OpenStudy (aravindg):